# Is Cognition Consistent with Perception? Assessing and Mitigating Multimodal Knowledge Conflicts in Document Understanding

### v2

@@ 5.2 C&P Consistency Results
The evaluation is conducted on the dataset constructed in Section 2.2. The experimental results, presented in Table 1, demonstrate that our multimodal knowledge consistency fine-tuning method enhances C&P consistency across all five datasets. Specifically, InternVL2-2b and InternVL2-8b show improvements of 3.72% and 3.60% in C&P consistency, respectively, while Qwen2.5-VL-7b exhibits a 3.36% increase. Under ideal conditions, consistency also improves. These findings indicate that our method effectively reduces C&P knowledge conflicts by linking perceptual and cognitive tasks. The comparison between Qwen2.5-VL-7b and the InternVL2 models highlights the general applicability of our approach across different MLLM architectures.
# Link. Conn. Doc VQA Deep Form Chart QA Average
1 80.59 90.62 72.05 77.40 83.80 91.27 76.40 86.05
2 82.97 91.52 77.85 80.97 87.45 93.66 79.10 87.77
3 82.71 91.14 77.24 80.72 87.45 93.26 79.24 87.51
4 83.39 91.32 78.56 82.20 87.64 93.17 80.12 88.21
@@ 5.3 The Performance of Cognitive and Perceptual Tasks
To assess the impact of C&P consistency on model performance, we evaluate the model’s effectiveness on cognitive and perceptual tasks. For the cognitive task, following previous works (Borchmann et al., 2021; Lee et al., 2023; Luo et al., 2024), we evaluate DocVQA and FUNSD using ANLS (Biten et al., 2019), DeepForm using the F1 score, and ChartQA using relaxed accuracy (Methani et al., 2020). For the perceptual task, all datasets are evaluated using ANLS.
As shown in Table 2, the three MLLMs show improved performance on both cognitive and perceptual tasks across all datasets after the multimodal knowledge consistency fine-tuning. We attribute this improvement to our fine-tuning approach, which reduces the conflict between perceptual and cognitive knowledge, thereby promoting their integration. We believe that the results suggest that enhancing C&P consistency can strengthen the capabilities of MLLMs.
@@ 5.5 Analysis of Conflict Patterns and Case Evidence
To further evaluate the effectiveness of multimodal knowledge consistency fine-tuning, we reuse the procedure described in Section 3 to analyze conflict patterns. Figure 4 (a) shows that, after fine-tuning, character-level errors (P1) and cognitive bias (P2) decrease significantly, making limited cognitive ability (P3) the dominant pattern. This shift supports our claim that heterogeneous VQA and OCR annotations are the primary sources of C&P knowledge conflicts and confirms that our method mitigates them effectively. The qualitative evidence in Figure 4 (b) illustrates these statistics. In case (b-1), categorized as P2, both InternVL2-2b and InternVL2-8b recognize “sustainable development” but incorrectly respond with “sustainability development” in the VQA task due to cognitive bias. The conflicts disappear after fine-tuning, as the models better integrate cognitive and perceptual knowledge. Notably, a similar P1 case with the same conclusion is provided in the Section A.7. In case (b-2), categorized as P3, the result indicates that our method cannot fundamentally extend the model’s cognitive boundaries.
@@ 6 Related Work
Document understanding (Cui et al., 2021; Xu et al., 2020, 2021; Huang et al., 2022; Luo et al., 2023, 2024; Wang et al., 2023) is a rapidly growing research area driven by increasing industrial demand. Its main objective is to comprehend complex typeset images that contain rich textual information, such as scanned document pages (Mathew et al., 2021; Svetlichnaya, 2020; Stanisławek et al., 2021), charts (Masry et al., 2022; Kafle et al., 2018; Methani et al., 2020), tables (Pasupat and Liang, 2015; Chen et al., 2019), and other formats (Tanaka et al., 2021; Mathew et al., 2022). As a multimodal task, document understanding involves automated processes for understanding, classifying, and extracting information, requiring models to possess both perceptual and cognitive capabilities (Cui et al., 2021). Recent studies (Chen et al., 2024; Hong et al., 2024; Dong et al., 2024; Bai et al., 2025) for general MLLMs improve the encoding resolution of document images, significantly boosting performance in document understanding tasks. Several MLLMs are developed to focus on addressing document understanding problems, such as mPLUG-DocOwl (Ye et al., 2023a; Hu et al., 2024) and UReader (Ye et al., 2023b).
@@
LLMs are distinguished for encapsulating an extensive repository of world knowledge, known as the memory. Simultaneously, LLMs continue to engage with external contextual knowledge post-deployment (Pan et al., 2023). The discrepancies between the contexts and the model’s memory knowledge, i.e. context-memory conflicts, are being intensively studied recently (Xie et al., 2023). Another notable challenge arises with intra-memory conflict—a condition where LLMs exhibit unpredictable behaviors to inputs that are semantically equivalent but syntactically distinct (Chang and Bergen, 2023; Rabinovich et al., 2023; Bartsch et al., 2023; Li et al., 2024b; Zhu et al., 2024; Zhang et al., 2024). This variance can be attributed to the conflicting knowledge embedded within the LLM’s memory, which stems from the inconsistencies present in the complex and diverse pre-training datasets.
@@
MLLMs provide powerful tools for content generation across a wide range of tasks. However, they are susceptible to hallucinations (Bang et al., 2023; Zhang et al., 2023b; Liu et al., 2024b), where the generated outputs contain information not present in the visual input. These hallucinations typically arise when the models overly rely on the strong priors of their language modules. Such conflicts between MLLMs’ language and visual perception raise concerns about their reliability and limit their applications (Ji et al., 2023; Kaddour et al., 2023). Current research primarily focuses on detecting and evaluating hallucinations (Li et al., 2023; Zhang et al., 2023b), as well as methods to reduce them (Liu et al., 2024a; Wang et al., 2024). To mitigate hallucinations, efforts have been directed toward enhancing data collection and training procedures (Liu et al., 2024a; Wang et al., 2024). Nevertheless, research on how MLLMs integrate perception and cognition knowledge, which is also vital for interpreting and debugging these models, has not progressed at the same pace.
@@ A.1 Details of Selected Datasets
Document QA. DocVQA (Mathew et al., 2021) contains 50k question-answer pairs from 12k document images in the UCSF Industry Documents Library. DUDE (Van Landeghem et al., 2023) covers diverse domains, including medical, legal, technical, and financial, providing 41k question-answer pairs from 5k documents. We exclude all multi-page VQA annotations from DUDE, retaining only single-page annotations.
Document IE. DeepForm (Svetlichnaya, 2020) and FUNSD (Jaume et al., 2019) are two Information Extraction datasets. DeepForm consists of 1.1k documents related to election spending. FUNSD contains 0.2k document images from the RVL-CDIP dataset (Harley et al., 2015). The annotations for DeepForm and FUNSD are transformed into a question-answer format, with DeepForm following Hu et al. (2024), and FUNSD following Luo et al. (2024). The annotations in Hu et al. (2024) for DeepForm incorrectly assume that all key values are on the first page, ignoring that DeepForm documents are multi-page. We correct this issue (see Section A.3 for details), ensuring information extraction occurs on the correct pages.
Chart QA. ChartQA (Masry et al., 2022) compiles a diverse range of topics and chart types from four primary sources: Statista (statista.com), The Pew Research Center (pewresearch.org), Our World in Data (ourworldindata.org), and the OECD (oecd.org). In total, the dataset includes 21k chart images and 32k question-answer pairs.
@@ A.2 Details of Evaluation Sample Construction
Doc VQA DUDE Deep Form FUNSD Chart QA
@@ A.3 Details of DeepForm Single-page QA Annotations
As described in Section A.1, Hu et al. (2024) provide incorrect annotations for DeepForm because they assume all key values are on the first page, overlooking that DeepForm documents are multi-page. To address this, we use GPT-4o to identify the correct page for information extraction using the prompt detailed in Table 12, ensuring that all single-page QA annotations in DeepForm are correct.
@@ A.4 Additional Details of C&P Knowledge Conflicts Evaluation
DocVQA DUDE DeepForm FUNSD ChartQA Average
GPT-4o 85.58 93.35 67.84 87.30 62.70 71.20 78.76 90.52 81.41 92.38 75.26 86.95
Qwen-VL-Max 95.66 97.15 82.54 91.35 83.23 86.69 83.19 90.52 92.44 95.68 87.41 92.28
@@
A.5 Additional Results of the Ablation Study
@@
A.6 Additional Details of C&P Connector
@@
A.7 Additional Case Study

### v3

@@ 5.2 C&P Consistency Results
The evaluation is conducted on the dataset constructed in Section 2.2. The experimental results, presented in Table 1, demonstrate that our multimodal knowledge consistency fine-tuning method enhances C&P consistency across all five datasets. Specifically, InternVL2-2b and InternVL2-8b show improvements of 3.72% and 3.60% in C&P consistency, respectively, while Qwen2.5-VL-7b exhibits a 3.36% increase. Under ideal conditions, consistency also improves. These findings indicate that our method effectively reduces C&P knowledge conflicts by linking perceptual and cognitive tasks. The comparison between Qwen2.5-VL-7b and the InternVL2 models highlights the general applicability of our approach across different MLLM architectures. Additionally, we perform two-sided paired t-tests using InternVL2-2b in Section A.9, showing that all gains in Table 1 are statistically significant.
# Link. Conn.
Doc
VQA
Deep
Form
Chart
QA
Average
1
80.59
90.62
72.05
77.40
83.80
91.27
76.40
86.05
2
82.97
91.52
77.85
80.97
87.45
93.66
79.10
87.77
3
82.71
91.14
77.24
80.72
87.45
93.26
79.24
87.51
4
83.39
91.32
78.56
82.20
87.64
93.17
80.12
88.21
@@ 5.3 The Performance of Cognitive and Perceptual Tasks
To assess the impact of C&P consistency on model performance, we evaluate the model’s effectiveness on cognitive and perceptual tasks. For the cognitive task, following previous works (borchmann2021due; lee2023pix2struct; layoutllm), we evaluate DocVQA and FUNSD using ANLS (biten2019scene), DeepForm using the F1 score, and ChartQA using relaxed accuracy (methani2020plotqa). For the perceptual task, all datasets are evaluated using ANLS.
As shown in Table 2, the three MLLMs show improved performance on both cognitive and perceptual tasks across all datasets after the multimodal knowledge consistency fine-tuning. We attribute this improvement to our fine-tuning approach, which reduces the conflict between perceptual and cognitive knowledge, thereby promoting their integration. We believe that the results suggest that enhancing C&P consistency can strengthen the capabilities of MLLMs. Similar to Section 5.2, the t-tests in Section A.9 show that the performance gains are statistically significant.
@@ 5.5 Analysis of Conflict Patterns and Case Evidence
To further evaluate the effectiveness of multimodal knowledge consistency fine-tuning, we reuse the procedure described in Section 3 to analyze conflict patterns. Figure 4 (a) shows that, after fine-tuning, character-level errors (P1) and cognitive bias (P2) decrease significantly, making limited cognitive ability (P3) the dominant pattern. This shift supports our claim that heterogeneous VQA and OCR annotations are the primary sources of C&P knowledge conflicts and confirms that our method mitigates them effectively. The qualitative evidence in Figure 4 (b) illustrates these statistics. In case (b-1), categorized as P2, both InternVL2-2b and InternVL2-8b recognize “sustainable development” but incorrectly respond with “sustainability development” in the VQA task due to cognitive bias. The conflicts disappear after fine-tuning, as the models better integrate cognitive and perceptual knowledge. Notably, a similar P1 case with the same conclusion is provided in the Section A.8. In case (b-2), categorized as P3, the result indicates that our method cannot fundamentally extend the model’s cognitive boundaries. In Figure 4, the responses of InternVL2-2b and InternVL2-8b are identical, reflecting the representativeness of these cases, though they differ in most other cases.
@@ 6 Related Work
Document understanding (cui2021document; xu2021layoutlmv2; huang2022layoutlmv3; luo2023geolayoutlm; layoutllm; shao2023gem; wang2023docllm; Zhu_2025_CVPR; mo2025doc) is a rapidly growing research area driven by increasing industrial demand. Its main objective is to comprehend complex typeset images that contain rich textual information, such as scanned document pages (docvqa; deepform; klc), charts (chartqa; kafle2018dvqa; methani2020plotqa), tables (wtq; chen2019tabfact; mo2025tablenarrator), and other formats (tanaka2021visualmrc; mathew2022infographicvqa; xing2024dochienet; shao2024webrpg). As a multimodal task, document understanding involves automated processes for understanding, classifying, and extracting information, requiring models to possess both perceptual and cognitive capabilities (cui2021document). Recent studies (internvl; hong2024cogagent; dong2024internlm; bai2025qwen2) for general MLLMs improve the encoding resolution of document images, significantly boosting performance in document understanding tasks. Several MLLMs are developed to focus on addressing document understanding problems, such as mPLUG-DocOwl (ye2023mplugdocowl; docowl15) and UReader (ye2023ureader).
@@
LLMs are distinguished for encapsulating an extensive repository of world knowledge, known as the memory. Simultaneously, LLMs continue to engage with external contextual knowledge post-deployment (pan2023knowledge). The discrepancies between the contexts and the model’s memory knowledge, i.e. context-memory conflicts, are being intensively studied recently (xie2023adaptive). Another notable challenge arises with intra-memory conflict—a condition where LLMs exhibit unpredictable behaviors to inputs that are semantically equivalent but syntactically distinct (chang2023languagemodelbehaviorcomprehensive; bartsch2023self; gvconsistency; zhu2024unraveling; zhang2024cross). This variance can be attributed to the conflicting knowledge embedded within the LLM’s memory, which stems from the inconsistencies present in the complex and diverse pre-training datasets.
@@
MLLMs provide powerful tools for content generation across a wide range of tasks. However, they are susceptible to hallucinations (bang-etal-2023-multitask; zhang2023sirenssongaiocean; liu2024insight), where the generated outputs contain information not present in the visual input. These hallucinations typically arise when the models overly rely on the strong priors of their language modules. Such conflicts between MLLMs’ language and visual perception raise concerns about their reliability and limit their applications (Ji_2023; kaddour2023challenges). Current research primarily focuses on detecting and evaluating hallucinations (li2023evaluating; zhang2023sirenssongaiocean), as well as methods to reduce them (liu2024mitigating; hevigc2024). To mitigate hallucinations, efforts have been directed toward enhancing data collection and training procedures (liu2024mitigating; hevigc2024). Nevertheless, research on how MLLMs integrate perception and cognition knowledge, which is also vital for interpreting and debugging these models, has not progressed at the same pace.
@@ 7 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 62372408).
@@ A.1 Details of Selected Datasets
Document QA. DocVQA (docvqa) contains 50k question-answer pairs from 12k document images in the UCSF Industry Documents Library. DUDE (dude) covers diverse domains, including medical, legal, technical, and financial, providing 41k question-answer pairs from 5k documents. We exclude all multi-page VQA annotations from DUDE, retaining only single-page annotations.
Document IE. DeepForm (deepform) and FUNSD (funsd) are two Information Extraction datasets. DeepForm consists of 1.1k documents related to election spending. FUNSD contains 0.2k document images from the RVL-CDIP dataset (harley2015evaluation). The annotations for DeepForm and FUNSD are transformed into a question-answer format, with DeepForm following docowl15, and FUNSD following layoutllm. The annotations in docowl15 for DeepForm incorrectly assume that all key values are on the first page, ignoring that DeepForm documents are multi-page. We correct this issue (see Section A.3 for details), ensuring information extraction occurs on the correct pages.
Chart QA. ChartQA (chartqa) compiles a diverse range of topics and chart types from four primary sources: Statista (statista.com), The Pew Research Center (pewresearch.org), Our World in Data (ourworldindata.org), and the OECD (oecd.org). In total, the dataset includes 21k chart images and 32k question-answer pairs.
@@ A.2 Details of Evaluation Sample Construction
Doc
VQA
DUDE
Deep
Form
FUNSD
Chart
QA
@@
DocVQA DUDE DeepForm FUNSD ChartQA Average
GPT-4o
85.58
93.35
67.84
87.30
62.70
71.20
78.76
90.52
81.41
92.38
75.26
86.95
Qwen-VL-Max
95.66
97.15
82.54
91.35
83.23
86.69
83.19
90.52
92.44
95.68
87.41
92.28
@@ A.3 Details of DeepForm Single-page QA Annotations
As described in Section A.1, docowl15 provide incorrect annotations for DeepForm because they assume all key values are on the first page, overlooking that DeepForm documents are multi-page. To address this, we use GPT-4o to identify the correct page for information extraction using the prompt detailed in Table 14, ensuring that all single-page QA annotations in DeepForm are correct.
@@ A.4 Additional Details of C&P Knowledge Conflicts Evaluation
A.5 Additional Results of the Synthetic Noise Experiment
The additional results of the synthetic noise experiment based on InternVL2-2b (Section 3) show in Table 10.
@@ A.5 Additional Results of the Synthetic Noise Experiment
A.6 Additional Results of the Ablation Study
@@
A.7 Additional Details of C&P Connector
@@
DocVQA DUDE DeepForm FUNSD ChartQA Average
0%
80.59
90.62
64.69
83.00
72.05
77.40
80.84
87.95
83.80
91.27
76.40
86.05
5%
79.12
90.97
64.09
82.70
71.36
76.31
79.74
87.67
83.81
90.12
75.62
85.55
10%
79.01
90.08
62.58
81.80
72.17
76.85
80.18
87.34
83.24
90.39
75.44
85.29
20%
77.92
89.93
61.61
81.17
70.24
75.38
77.53
85.45
82.08
90.52
73.88
84.49
Table 10: The synthetic noise experiment based on InternVL2-2b. All values are percentages (%), where the primary number represents C&P Consistency and the smaller number represents Idealized C&P Consistency.
A.8 Additional Case Study
@@
A.9 Additional Results of the T-test
We perform two-sided paired t-tests using InternVL2-2b, and the results are shown in Table 11.
DocVQA DUDE DeepForm FUNSD ChartQA
C&P Consistency
5.22
[…]
4.75
[…]
5.41
[…]
2.36
[…]
4.72
[…]
Cognitive Task
4.69
[…]
3.28
[…]
3.32
[…]
2.10
[…]
4.55
[…]
Perceptual Task
6.52
[…]
4.49
[…]
5.81
[…]
2.53
[…]
3.20
[…]
Table 11: Results of two-sided paired t-tests using InternVL2-2b, reported as t-statistics with p-values in parentheses.